OpenStudy (anonymous):

q= (5q - 3n)/m = f i have no clue how to do this

OpenStudy (solomonzelman):

which variable are you solving for, q ?

OpenStudy (anonymous):

solving for q

OpenStudy (solomonzelman):

are you sure that the equation has to have 2 equal since, maybe one of them is a plus ?

OpenStudy (solomonzelman):

2 equal signs, not since ...

OpenStudy (anonymous):

q = (mf - 5)/3 q = (mf/5) + 3n q = m((5f +n)/3) q = (mf+3n)/5 this is my list of possible answers

OpenStudy (solomonzelman):

your equation should be fixed.

OpenStudy (anonymous):

Elaborate

geerky42 (geerky42):

In your question, you said q= (5q - 3n)/m = f should that second equal sign supposed to be +?

OpenStudy (solomonzelman):

Well, look at what you wrote. I think you wrote the equation in which you are to solve for q incorrectly.

OpenStudy (anonymous):

it won't come out as a fraction

OpenStudy (solomonzelman):

(m+f) is the denominator ?

OpenStudy (anonymous):

5q-3n is numerator and m is denominator = f

OpenStudy (anonymous):

solving for q

geerky42 (geerky42):

Is your question read as \(\Large q= \dfrac{5q - 3n}{m} = f\) ?

OpenStudy (anonymous):

yes